OpenStudy (anonymous):

Find the partial derivatives indicated. Assume the variables are restricted to a domain on which the function is defined. z=(x4+x−y)^4 ∂z∂x= ∂z∂y= Find the partial derivatives indicated. Assume the variables are restricted to a domain on which the function is defined. z=(x4+x−y)^4 ∂z∂x= ∂z∂y= @Mathematics

OpenStudy (anonymous):

dz/dx = 4(x^4 + x -y)^3(4x^3+1) dz/dy = -4((x^4 +x -y)^3